# Impact of boiling liquid droplets: Vapor entrapment suppression

### v1

@@ Conclusions
We create an approximate model based on scaling arguments to explain our findings, which is in good agreement with our experiments. In addition, we devised a numerical simulation consisting of a boundary integral method for the impacting droplet, coupled to a lubrication model including phase change for the vapor layer that provides insight into the vapor dynamics below the impacting droplet.
Our research is of consequence for the transport of cryogenic liquids such as LNG and LH2, where the suppression of vapor bubble entrapment is likely to go hand in hand with increased impact pressures. The models introduced in this work may be translated to the conditions in those applications to verify under which circumstances phase change may become a non-negligible effect and needs to be included into the engineering of cryogenic liquid carriers on an industrial scale.

### v2

@@ Conclusions
We create an approximate model based on scaling arguments to explain our findings, which is in good agreement with our experiments. The argument starts from the same mass-energy balance approximation at the liquid-vapor interface we also used successfully in explaining the strongly increased pressures observed during wave [4] and disc [5] impact in a boiling liquid, which provides additional confidence in the correctness of this approach.
In addition, we devised a numerical simulation consisting of a boundary integral method for the impacting droplet, coupled to a compressible lubrication model including phase change for the vapor layer that provides insight into the vapor dynamics below the impacting droplet. Since gas compressibility effects in droplet impact are important and constitute an active research area [9, 28], it is good to stress that we are in an impact regime where compressibility effects are expected to be minor [21], as discussed in the SI. However, phase change (condensation) can only be reliably incorporated into a compressible lubrication model.
Our research is of consequence for the transport of cryogenic liquids such as LNG and LH2, where the suppression of vapor bubble entrapment is likely to go hand in hand with increased impact pressures. The models introduced in this work may be translated to the conditions in those applications to verify under which circumstances phase change may become a non-negligible effect and needs to be included into the engineering of cryogenic liquid carriers on an industrial scale. More specifically, using our theoretical results in estimating when the rate of condensation exceeds the impact speed of a sloshing wave crest, will make it possible to identify and avoid situations in which the protective cushioning effect of the intermediate vapor phase is cancelled by condensation. The latter will increase the impact load, potentially threatening the structural integrity of the transporting containers.